### v1

# Experimental Multi-site Testbed for Advanced Control and Optimization of Hybrid Energy Systems

Abstract
This paper presents a hybrid energy system (HES) experimental testbed developed at the University of Vermont to support prototyping and validation of advanced control and optimization strategies for grid services. The platform integrates hardware-in-the-loop (HIL) simulation with a reconfigurable set of kilowatt-scale assets, including solar photovoltaic (PV), battery storage, an electrolyzer as a controllable load, and grid-tied inverters. A unified monitoring and communication architecture supports real-time data acquisition, model validation, and control implementation. The testbed’s capabilities are demonstrated through a controller hardware-in-the-loop (CHIL) experiment in which a battery system participates in PV power smoothing.
I Introduction
The global transition toward sustainable energy systems has accelerated the deployment of renewable generation. However, the inherent variability and uncertainty of renewable resources pose significant challenges to grid reliability, flexibility, and power quality [liang2016emerging]. Hybrid Energy Systems (HES) have emerged as a promising solution to address the growing global demand for reliable, efficient, and sustainable power [hassan2023review]. HES integrate diverse energy resources, such as intermittent renewable sources like photovoltaic (PV) and wind generation, energy storage, and controllable loads including electrolyzers, electric vehicle (EV) charging stations, or data centers to enhance flexibility and reliability compared to standalone systems [nassar2025role]. This has driven the rapid expansion of HES installations in the United States as shown in Fig. 1 [Berkeley2025].
Deploying HES at scale introduces new layers of operational and control complexity, including power converter interactions, energy and ramping constraints, coupled dynamics, and multi-timescale coordination [das2025research]. Most existing research is either limited to pairs of HES components (e.g., PV+X+X) or relies on mathematical modeling and digital simulation. Consequently, these studies are primarily evaluated based on model accuracy, computational efficiency, and the convergence of numerical solvers. Furthermore, these models often fail to capture nonlinear switching effects, communication delays, hardware limits, and electromagnetic compatibility issues. Conversely, full-scale field testing is often infeasible due to cost, safety concerns, and complexity. This creates a validation gap, highlighting the need for intermediate methodologies combining hardware testing fidelity with laboratory controllability and safety [stanev2020real].
Figure 1: Operational HES across the U.S. [Berkeley2025].
TABLE I: Summary of Selected North America testbeds and their key assets and capabilities
Testbed Location Scale Key Assets Capabilities HES Integration Dual-site
Energy System Integration Facility (NREL) [kroposki2012energy] Colorado 1 MW Power converters, batteries, electrolyzer, and fuel cell HIL validation, power management evaluation, power system emulation, and microgrid control. Yes Yes
HIL Platform (North Carolina State University) [ghanbari2019hardware] North Carolina 250 kW PV and power converters MPPT implementation, voltage and frequency control, grid emulation, and HIL validation. No No
ALSET Lab (Rensselaer Polytechnic Institute) [chang2024power] New York 150 kW Inverter and power supply PV emulation, HIL validation, Volt-VAR/Watt modes, and low voltage ride-through tests. No No
FIU testbed (Florida International University) [youssef2015dds] Florida 40 kW AC generators, power converters, adjustable loads, and SCADA system PV and wind emulation, HES AC/DC control, energy management, and MPPT implementation. Limited No
SSMTB testbed (Sandia) [glover2012secure] New Mexico 30 kW Diesel generators, energy storage, and programmable loads PV and wind emulation, voltage and frequency control, and HIL validation. Limited No
UTA testbed (University of Texas-Arlington) [turner2014design] Texas 15 kW PV, wind turbine, fuel cell, battery, power converters, and loads Power management, voltage and frequency control, and power quality analysis. Yes No
Laboratory-based testbeds bridge this validation gap by providing controlled and repeatable environments to study complex device interactions. Hardware-in-the-loop (HIL) simulation in particular has become critical for these validation efforts. In an HIL configurations, physical hardware components, such as digital controllers or power electronic devices, are interfaced with a real-time digital simulation of a power system network. This approach enables the systematic evaluation of device-level performance, controller robustness, and overall system-level stability before field deployment. Several experimental HIL testbeds have been developed, addressing different aspects of HES validation and inverter-dominated grid integration, as summarized in Table I. While these facilities demonstrate valuable capabilities, most focus on either large-scale grid integration or individual component validation. However, there remains a critical need for reconfigurable, intermediate-scale platforms that integrate multiple HES resources, including electrolyzers, and are explicitly designed to study coordinated multi-component HES control and co-optimization across multiple timescales.
This paper presents a new HES experimental testbed at the University of Vermont (The Accelerated Testing Lab), designed for the prototyping and validation of advanced control and coordination strategies to provide grid services. Compared to existing testbeds (Table I), the platform uniquely combines:
  • A dual-site architecture with field PV and weather data driving HIL experiments,
  • A reconfigurable HES platform integrating PV, battery, inverters, and electrolyzer with plug-and-play interfaces,
  • A unified framework for multi-timescale HES control.
This paper provides a comprehensive overview of the key capabilities and features of this testbed, and experimentally validates battery participation in solar smoothing as a representative test case.
The remainder of this paper is organized as follows. Section II describes the HES laboratory design and HIL configuration of the testbed, including the electrical system architecture and data streaming framework. Section III details the key capabilities of the HES platform. Section IV presents an experimental test case, while Section V concludes the paper with perspectives on ongoing and future work.
II Experimental testbed and system architecture
The novel HES testbed features a distributed architecture composed of two physically distinct but tightly interconnected laboratory facilities. These facilities share real-time data, control signals, and HIL interfaces, enabling coordinated experiments that span devices, controllers, and grid-level scenarios.
  • The Accelerated Testing Laboratory (ATL), the primary on-campus HIL platform.
  • The Hybrid Solar Test Center (HSTC), an off-campus 23 kW solar and meteorological site, where a 100kW/100kWh battery storage will be installed in 2026.
This two-part structure enables model validation, rapid control prototyping, and HIL analysis at the ATL using real-world solar and weather data from HSTC. The overall system architecture is shown in Fig. 2. The specific ratings for all key assets are listed in Table II. The design and capabilities of each facility are detailed in the following subsections.
Figure 2: HES testbed showing ATL components, including inverters interfaced with the RGS on the AC side, battery modules powered through the DC supply, an electrolyzer as a controllable load, and OPAL-RT integration for HIL testing; and HSTC components including the PV array and planned battery system.
TABLE II: Equipment Specifications
Component Description Specifications
PV Array ET Solar 23.4 kW DC
Inverter SolarEdge 3×10 kVA, single-phase
Battery (TBI 2026) KORE Power 100 kW/100 kWh
LFP Battery Enphase ACB 1.0 6×1.2 kWh, 26.5V
NMC Battery KORE Power 2×6.5 kWh, 59.6V
Inverter UNIFI reference design 4×5 kVA, 3-phase
Electrolyzer Verde hydrogen 15 kW alkaline stack
DC Power Supply Chroma 62120D-1200 12 kW, 0–1200V, $\pm$55A
Regenerative Grid Simulator Chroma 61812 12 kVA, 0–600V
Microcontroller TI C2000 100 MHz
Real-Time Simulator Opal-RT OP5033 Intel® Xeon® (4 cores)
FPGA and I/O Expansion Opal-RT OP4610 FPGA, 64 A/D I/O
II-A Accelerated Testing Laboratory
The ATL at the University of Vermont (Fig. 2) integrates hardware and simulation capabilities to enable comprehensive testing of power electronics, energy storage, and control strategies in a reconfigurable cyber-physical environment. The bidirectional DC power supply feeds either the DC bus of two 5 kVA inverters connected in parallel or the batteries. The inverters, designed based on the UNIFI reference design [UNIFI2025], enable testing advanced control strategies and detailed investigation of inverter dynamics and their impact on the grid. Each inverter is controlled by a Texas Instruments (TI) TMDSCNCD28379D microcontroller, which enables the implementation of advanced control strategies by sensing analog voltage and current signals through ADC cards and generating required PWM signals based on control modes such as grid-forming (GFM) and grid-following (GFL) control, active/reactive power regulation, and coordinated multi-inverter operation. In addition, the DC power supply is connected to six 1.2 kWh Enphase LFP batteries, allowing for systematic charge, discharge, and storage co-optimization studies. On the AC side, the inverters are connected to a regenerative grid simulator (RGS). The testbed is coupled to a 15 kW electrolyzer stack, which represents hydrogen production as a controllable load interfaced with the grid.
Furthermore, the Opal-RT simulator, with the OP4610 I/O expansion unit, provides the computational environment to emulate detailed grid behaviors. As shown in Fig. 2, the testbed architecture is explicitly designed for ”plug-and-play” flexibility. This allows any component initially simulated within the Opal-RT to be seamlessly replaced by its physical hardware counterpart for full HIL validation. This creates a robust cyber-physical loop where Opal-RT issues analog setpoints to the RGS and bidirectional DC supply, which act as the physical interface to the device under test. These controllable power supplies generate the precise power and voltage conditions dictated by the real-time simulation. Voltage and current sensors capture responses from the hardware and feed them back to the Opal-RT’s analog inputs, completing the HIL setup. This closed-loop configuration ensures that the real-time model dynamically reacts to physical hardware behavior and vice versa. This setup also facilitates integration of PV and weather profiles from the HSTC (discussed in the next subsection) to drive realistic validation scenarios.
II-B Hybrid Solar Test Center
The ATL is supported by the HSTC located at the McNeil generating plant roughly one mile from campus. The HSTC hosts a 23 kW grid-connected solar PV array interfaced through SolarEdge inverters, along with an on-site weather station for real-time meteorological data acquisition. The weather station records solar irradiance, temperature, wind speed, and humidity every 5 seconds, providing synchronized environmental and electrical data for PV modeling and performance analysis. Future expansion includes the installation of a 100 kW/100 kWh battery energy storage system (BESS) in 2026, enabling field validation of battery-solar coordination strategies and grid services under real-world conditions.
Figure 3: Data and communication architecture between HSTC and ATL.
Environmental and operational data are collected from local and remote sources, as shown in Fig. 3, to support real-time monitoring, control, and performance evaluation of all HES subsystems within the ATL. At the inverter and PV panel level, data from HSTC including AC and DC voltage, current, power, and frequency are measured using high-precision sensors and transmitted to the server through Modbus energy meter coupled with a communication gateway via Starlink network. All data streams are time-synchronized and stored on a dedicated server for visualization, analysis, and archival purposes.
The setup provides a unique and unified platform for validating HES control strategies and system models, as well as assessing system dynamic behavior and grid-integration performance under controlled conditions. The key capabilities of the testbed are discussed in the following section.
III Platform Capabilities and Features
The experimental testbed provides a comprehensive environment for investigating HES integration and advanced control validation. The platform key capabilities span three interconnected domains:
  • Component and system-level modeling and validation
  • Control implementation and verification
  • Hardware-in-the-loop (HIL) integration
This section describes the platform capabilities and their application to contemporary challenges in power system operation.
III-A Component and system-level modeling and validation
Accurate models of individual HES components and their interactions are fundamental to enabling informed decision-making in system design, control development, and operational optimization. The ATL provides comprehensive experimental capabilities to validate physics-based and data-driven models across multiple components and timescales. The bidirectional DC power supply enables experimental validation of battery models through controlled charge-discharge cycling, including the energy reservoir model [omidi2025towards] and electro-thermal equivalent circuit model [mishra2024predictive] developed in prior work. For the electrolyzer subsystem, physics-based models describing hydrogen production rate, efficiency, and thermal dynamics have been developed and experimentally validated through systematic power cycling experiments on the 15 kW alkaline stack [ayubirad2025modeling].
The platform further enables validation of efficiency-based and detailed switching inverter models. Data from the HSTC support validation of PV models across diverse operating conditions and temperature dependencies. At the system level, aggregated circuit models capture interactions between PV, battery, and inverter systems through equivalent circuit modeling frameworks, facilitating consistent state estimation across component boundaries and robust parameter identification [sang2024circuit]. These validated models form the foundation for multi-objective optimization and control studies.
III-B Control implementation and verification
HES change how controllers must be designed. While individual components can often be controlled independently, the strong electrical and thermal coupling in HES makes coordinated controller design both necessary and beneficial. As power systems become increasingly inverter dominated, traditional sources of inertia and voltage support disappear, and these services must be provided through tightly integrated, multi-layer control across HES components. Addressing these challenges requires experimental validation of advanced control and coordination strategies under realistic operating conditions. The ATL testbed bridges this gap by providing a flexible platform to validate the control algorithms at both the component and system levels. At the component level, ATL allows direct implementation and testing of inverter and battery control algorithms, such as GFM and GFL modes, active/reactive power regulation, Volt-VAR control, and multi-inverter coordination, under programmable conditions including voltage sags, weak grid operation, and fault events. At the system level, the platform supports hierarchical and distributed control schemes, including model predictive control and economic dispatch-based coordination of batteries, PV, and electrolyzers. These validated control strategies enable the ATL to test algorithms for providing different grid services, summarized in Table III. The shaded grid service is experimentally validated in the next section.
TABLE III: Grid Services Capabilities of Campus Testbed
Grid Services Time scale Participating assets
Fast Frequency Response <<1s Battery+Inverter
Frequency Regulation seconds-minutes Battery/Coupled HES
[…]
Volt-Var and voltage regulation 100ms-10s Battery/PV + Inverter
Power Quality <<1s Inverter
Energy shifting and peak shaving minutes-hours Coupled HES
III-C Hardware-in-the-loop (HIL) integration
HIL validation bridges the gap between simulation and field deployment by integrating real hardware with real-time digital simulation. It enables realistic evaluation of controllers, converters, and system interactions while maintaining safety and flexibility. The ATL testbed, equipped with an OPAL-RT real-time simulator and I/O unit, supports both CHIL and HIL. The OPAL-RT simulator supports steady-state, transient, and dynamic grid models, enabling comprehensive evaluation of hardware interactions with the grid across different timescales. This enables detailed studies of inverter–grid interactions, coordination among HES components, and validation of advanced control strategies under realistic operating scenarios.
IV Experimental Test Case: Solar Smoothing using Physical Battery Storage
The increasing penetration of PV generation introduces rapid power fluctuations from cloud transitions and irradiance variability, which challenges grid stability and power quality. A BESS can mitigate these fluctuations through fast response and bidirectional power capability, absorbing excess power during PV spikes and injecting power during drops. By coordinating PV generation with BESS operation, HES can smooth the power delivered to the grid. Grid operators impose ramp-rate limits to keep fluctuations within acceptable ranges, defined as:
$\text{Ramp Rate: }RR(t)=\frac{P_{\text{PV}}(t)-P_{\text{PV}}(t-\Delta t)}{\Delta t\cdot P_{\text{rated}}}\times 100$ (1)
where $P_{\text{PV}}(t)$ is the PV power at time $t$, $\Delta t$ is the time interval, and $P_{\text{rated}}$ is the rated PV capacity. In this study, the maximum allowable ramp rate is set to $\pm 5\%$/min, which lies within the 2-10%/min range specified in several PV grid codes studies.
To evaluate the smoothing capability at ATL, a CHIL experiment was conducted using the experimental setup shown in Fig. 4. The experiment uses historical 5-second PV data from one inverter at the HSTC site. The PV profile is scaled to match the available battery capacity; this scaling does not affect ramp-rate behavior because ramp rates are expressed relative to the rated capacity. The OPAL-RT sends PV power and measured battery voltage to the TI controller, which runs a 30-minute moving-average smoothing algorithm, detailed in Algorithm 1. The controller issues battery current setpoints to the power supply through OPAL-RT. The battery management system (BMS) estimates the battery state of charge (SOC).
Figure 4: Experimental setup for solar smoothing CHIL validation showing two series-connected Enphase batteries with a BMS, bidirectional DC power supply, TI control card, and OPAL-RT real-time simulator.
In : $P_{\mathrm{PV}}[k]$, $V_{\mathrm{batt}}[k]$
Out : $I_{\mathrm{set}}[k]$
// Initialization
$P_{\mathrm{buf}}[i]\leftarrow 0,\ i=1{:}N$; $k\leftarrow 1$
while controller is running do
  // 1) Insert new PV sample
  $P_{\mathrm{buf}}[([k-1]\bmod N)+1]\leftarrow P_{\mathrm{PV}}[k]$
  // 2) Compute smoothed power
  $\hat{P}_{\mathrm{PV}}[k]\leftarrow\frac{1}{N}\sum_{i=1}^{N}P_{\mathrm{buf}}[i]$
  // 3) Battery power
  $P_{\mathrm{batt}}[k]\leftarrow P_{\mathrm{PV}}[k]-\hat{P}_{\mathrm{PV}}[k]$
  // 4) Current setpoint
  $I_{\mathrm{set}}[k]\leftarrow\dfrac{P_{\mathrm{batt}}[k]}{V_{\mathrm{batt}}[k]}$
  // 5) Output and increment
  Output $I_{\mathrm{set}}[k]$
  $k\leftarrow k+1$ ; sleep($T_{s}$)
Algorithm 1 PV Smoothing Controller
A two-hour snapshot of the CHIL experiment, shown in Fig. 5, highlights the system smoothing performance. Fig. 5(a) shows the raw PV power and compares the smoothed results from simulation and experiment, which indicate that the results align closely. Fig. 5(b) presents the dispatched battery power, and Fig. 5(c) shows the resulting SOC, which remain within safe operating limits throughout the experiment. The raw PV profile shows a maximum ramp rate of 56%/min, exceeding the $\pm 5\%$/min grid limit, whereas the BESS effectively reduces the maximum ramp rate to 3.75%/min. The corresponding distribution of ramp rates is shown in Fig. 5(d), which illustrates the significant reduction in extreme events. These results validate the capability of the ATL facility to effectively implement and test PV smoothing strategies using BESS.
Figure 5: Two-hour snapshot of HIL experimental PV power smoothing using historical 5-second HSTC data and ATL battery setup. (a) raw and smoothed PV power, here, $P_{\text{PV}}$ denotes the raw PV power, $\hat{P}_{\text{PV}}^{e}$ and $\hat{P}_{\text{PV}}^{s}$ denote the smoothed PV power from experiment and simulation, respectively, (b) dispatched battery power, (c) battery SOC, and (d) distribution of ramp rates.
V Conclusion
This paper presented an experimental testbed for validating HES coordination and control strategies. The ATL platform integrates kW-scale PV, batteries, grid-tied inverters, and an electrolyzer with a real-time simulator. It enables modeling, control, and HIL validation of HES at both component and system levels. The solar smoothing experiment demonstrates ATL’s ability to test grid-service control strategies.
Future work will use the OPAL-RT real-time simulator to coordinate all subsystems within a unified control and communication framework targeting diverse operational objectives. Another direction will focus on implementing and validating GFL and GFM control strategies under weak-grid and transient conditions. These advancements will further establish the ATL as a comprehensive real-time facility for accelerated research and deployment of next-generation HES.
Acknowledgment
The authors thank Peng Sang and Madiha Akbar at the University of Vermont for contributions to PV data streaming and to modeling and testing the electrolyzer in the ATL.

### v2

@@
-Experimental Multi-site Testbed for Advanced Control and Optimization of Hybrid Energy Systems
+arXiv:2511.16819v2[eess.SY] 26 Feb 2026
 Abstract
-This paper presents a hybrid energy system (HES) experimental testbed developed at the University of Vermont to support prototyping and validation of advanced control and optimization strategies for grid services.
-The platform integrates hardware-in-the-loop (HIL) simulation with a reconfigurable set of kilowatt-scale assets, including solar photovoltaic (PV), battery storage, an electrolyzer as a controllable load, and grid-tied inverters.
+This paper presents a hybrid energy system (HES) experimental testbed developed at the University of Vermont, featuring a dual-site architecture that integrates on-campus laboratory facility with an off-campus solar and meteorological station.
+This supports the prototyping and validation of advanced HES control and optimization strategies.
+The platform integrates hardware-in-the-loop (HIL) simulations with a reconfigurable set of kVA-scale assets.
 A unified monitoring and communication architecture supports real-time data acquisition, model validation, and control implementation.
-The testbed’s capabilities are demonstrated through a controller hardware-in-the-loop (CHIL) experiment in which a battery system participates in PV power smoothing.
+The capabilities of the testbed are demonstrated through an HIL experiment in which a battery systems participate in solar PV smoothing.
 I Introduction
-The global transition toward sustainable energy systems has accelerated the deployment of renewable generation.
-However, the inherent variability and uncertainty of renewable resources pose significant challenges to grid reliability, flexibility, and power quality [liang2016emerging].
-Hybrid Energy Systems (HES) have emerged as a promising solution to address the growing global demand for reliable, efficient, and sustainable power [hassan2023review].
-HES integrate diverse energy resources, such as intermittent renewable sources like photovoltaic (PV) and wind generation, energy storage, and controllable loads including electrolyzers, electric vehicle (EV) charging stations, or data centers to enhance flexibility and reliability compared to standalone systems [nassar2025role].
-This has driven the rapid expansion of HES installations in the United States as shown in Fig.
-1 [Berkeley2025].
-Deploying HES at scale introduces new layers of operational and control complexity, including power converter interactions, energy and ramping constraints, coupled dynamics, and multi-timescale coordination [das2025research].
-Most existing research is either limited to pairs of HES components (e.g., PV+X+X) or relies on mathematical modeling and digital simulation.
-Consequently, these studies are primarily evaluated based on model accuracy, computational efficiency, and the convergence of numerical solvers.
-Furthermore, these models often fail to capture nonlinear switching effects, communication delays, hardware limits, and electromagnetic compatibility issues.
-Conversely, full-scale field testing is often infeasible due to cost, safety concerns, and complexity.
-This creates a validation gap, highlighting the need for intermediate methodologies combining hardware testing fidelity with laboratory controllability and safety [stanev2020real].
-Figure 1: Operational HES across the U.S. [Berkeley2025].
+The global transition toward sustainable energy has accelerated renewable energy deployment.
+However, the inherent variability and uncertainty of these resources challenges grid reliability, flexibility, and power quality [8].
+Hybrid Energy Systems (HES) have emerged as a promising solution to address the growing global demand for reliable, efficient, and sustainable power.
+HES integrate diverse energy resources, such as intermittent renewable sources like photovoltaic (PV) and wind generation, energy storage, and controllable loads including electrolyzers, electric vehicle (EV) charging stations, or data centers to enhance flexibility and reliability compared to standalone systems [11].
+This has driven the rapid expansion of HES installations in the United States  [16].
+Deploying HES at scale introduces new layers of operational and control complexity, including power converter interactions, energy and ramping constraints, coupled dynamics, and multi-timescale coordination [3].
+Most existing research is either limited to pairs of HES components (e.g., PV+X+X) or relies on mathematical modeling and simulation.
+Consequently, these studies are primarily evaluated based on model accuracy, computational efficiency, and numerical solvers.
+Furthermore, these models often fail to capture nonlinear switching effects, communication delays, and hardware limits.
+Conversely, full-scale field testing is often infeasible due to cost, safety issues, and complexity.
+This creates a validation gap, highlighting the need for intermediate methodologies combining hardware testing fidelity with laboratory controllability and safety [14].
+Laboratory-based testbeds bridge this validation gap by providing controlled and repeatable environments to study complex device interactions.
+Hardware-in-the-loop (HIL) simulation has become critical for these validation efforts.
+In an HIL configuration, physical hardware components are interfaced with real-time digital simulations to enable systematic evaluation of device and controller performance and system-level stability before field deployment.
+A multi-site HIL testbed refers to geographically separated facilities that exchange measurements and/or setpoints to support coordinated experiments.
+Such architectures can combine laboratory controllability with field realism, however they also introduce additional communication and integration requirements compared to single-site setups [4].
+Several HIL testbeds have been developed that address different aspects of HES validation and inverter-dominated grid integration, as summarized in Table I.
+While these facilities demonstrate valuable capabilities, most platforms emphasize either large-scale integration or single-facility component validation.
+Therefore, there remains a need for reconfigurable, intermediate-scale platforms that are multi-site and can incorporate field measurements into laboratory HIL studies, enabling multi-timescale HES control validation under real-world conditions.
 TABLE I: Summary of selected North America testbeds and their key assets and capabilities
-Testbed Location Scale Key Assets Capabilities HES Integration Dual-site
-Energy System Integration Facility (NREL) [kroposki2012energy] Colorado 1 MW Power converters, batteries, electrolyzer, and fuel cell HIL validation, power management evaluation, power system emulation, and microgrid control.
+Testbed Location Scale Key Assets Capabilities HES Integration Multi-site
+Energy System Integration Facility (NREL) [7] Colorado 1 MW Power converters, batteries, electrolyzer, and fuel cell HIL validation, power management evaluation, power system emulation, and microgrid control.
 Yes Yes
-HIL Platform (North Carolina State University) [ghanbari2019hardware] North Carolina 250 kW PV and power converters MPPT implementation, voltage and frequency control, grid emulation, and HIL validation.
-No No
-ALSET Lab (Rensselaer Polytechnic Institute) [chang2024power] New York 150 kW Inverter and power supply PV emulation, HIL validation, Volt-VAR/Watt modes, and low voltage ride-through tests. No No
-FIU testbed (Florida International University) [youssef2015dds] Florida 40 kW AC generators, power converters, adjustable loads, and SCADA system PV and wind emulation, HES AC/DC control, energy management, and MPPT implementation.
+HIL Platform (North Carolina State University) [5] North Carolina 250 kW PV and power converters MPPT implementation, voltage and frequency control, grid emulation, and HIL validation. No No
+FIU testbed (Florida International University) [17] Florida 40 kW AC generators, power converters, adjustable loads, and SCADA system PV and wind emulation, HES AC/DC control, energy management, and MPPT implementation.
 Limited No
-SSMTB testbed (Sandia) [glover2012secure] New Mexico 30 kW Diesel generators, energy storage, and programmable loads PV and wind emulation, voltage and frequency control, and HIL validation.
-Limited No
-UTA testbed (University of Texas-Arlington) [turner2014design] Texas 15 kW PV, wind turbine, fuel cell, battery, power converters, and loads Power management, voltage and frequency control, and power quality analysis.
+SSMTB testbed (Sandia) [6] New Mexico 30 kW Diesel generators, energy storage, and programmable loads PV and wind emulation, voltage and frequency control, and HIL validation. Limited No
+UTA testbed (University of Texas-Arlington) [15] Texas 15 kW PV, wind turbine, fuel cell, battery, power converters, and loads Power management, voltage and frequency control, and power quality analysis.
 Yes No
-Laboratory-based testbeds bridge this validation gap by providing controlled and repeatable environments to study complex device interactions.
-Hardware-in-the-loop (HIL) simulation in particular has become critical for these validation efforts.
-In an HIL configurations, physical hardware components, such as digital controllers or power electronic devices, are interfaced with a real-time digital simulation of a power system network.
-This approach enables the systematic evaluation of device-level performance, controller robustness, and overall system-level stability before field deployment.
-Several experimental HIL testbeds have been developed, addressing different aspects of HES validation and inverter-dominated grid integration, as summarized in Table I.
-While these facilities demonstrate valuable capabilities, most focus on either large-scale grid integration or individual component validation.
-However, there remains a critical need for reconfigurable, intermediate-scale platforms that integrate multiple HES resources, including electrolyzers, and are explicitly designed to study coordinated multi-component HES control and co-optimization across multiple timescales.
-This paper presents a new HES experimental testbed at the University of Vermont (The Accelerated Testing Lab), designed for the prototyping and validation of advanced control and coordination strategies to provide grid services.
+Multi-site networked HIL platform [4] Three states (CO,NC,TN) N/A ADMS, protocol adapters, HIL sites DER emulation, multi-site HIL evaluation of distributed control, and ADMS–DER coordination No Yes
+This paper presents a new multi-site HES experimental testbed at the University of Vermont (The Accelerated Testing Lab), designed for the prototyping and validation of advanced control and coordination strategies for grid services.
 Compared to existing testbeds (Table I), the platform uniquely combines:
-  • A dual-site architecture with field PV and weather data driving HIL experiments,
-  • A reconfigurable HES platform integrating PV, battery, inverters, and electrolyzer with plug-and-play interfaces,
-  • A unified framework for multi-timescale HES control.
+  • A dual-site architecture integrating off-site field PV and weather data with on-campus hardware assets and HIL,
+  • A reconfigurable HES platform with PV, batteries, inverters, and electrolyzer with plug-and-play interfaces,
+  • A unified framework for multi-timescale HES control and real-time data streaming.
 This paper provides a comprehensive overview of the key capabilities and features of this testbed, and experimentally validates battery participation in solar smoothing as a representative test case.
+The validation is conducted through an HIL setup with a physical battery, simulated batteries, and a real-time controller.
 The remainder of this paper is organized as follows.
@@
 Section III details the key capabilities of the HES platform.
-Section IV presents an experimental test case, while Section V concludes the paper with perspectives on ongoing and future work.
+Section IV presents an experimental test case, while Section V concludes the paper with perspectives on future work.
 II Experimental testbed and system architecture
-The novel HES testbed features a distributed architecture composed of two physically distinct but tightly interconnected laboratory facilities.
+The HES testbed features a distributed architecture composed of two physically distinct but tightly interconnected laboratory facilities:
+  • The Accelerated Testing Laboratory (ATL), the primary on-campus HIL platform.
+  • The Hybrid Solar Test Center (HSTC), an off-campus 23 kW solar and meteorological site with a planned battery system installation.
 These facilities share real-time data, control signals, and HIL interfaces, enabling coordinated experiments that span devices, controllers, and grid-level scenarios.
-  • The Accelerated Testing Laboratory (ATL), the primary on-campus HIL platform.
-  • The Hybrid Solar Test Center (HSTC), an off-campus 23 kW solar and meteorological site, where a 100kW/100kWh battery storage will be installed in 2026.
-This two-part structure enables model validation, rapid control prototyping, and HIL analysis at the ATL using real-world solar and weather data from HSTC.
 The overall system architecture is shown in Fig.
-2.
-The specific ratings for all key assets are listed in Table II.
+1, with key asset specifications in Table II.
 The design and capabilities of each facility are detailed in the following subsections.
-Figure 2: HES testbed showing ATL components, including inverters interfaced with the RGS on the AC side, battery modules powered through the DC supply, an electrolyzer as a controllable load, and OPAL-RT integration for HIL testing; and HSTC components including the PV array and planned battery system.
+Figure 1: Multi-site HES testbed with plug-and-play architecture.
+Components simulated in OPAL-RT can be replaced with physical ATL hardware (inverters, batteries, electrolyzer) or real-time HSTC field data (PV, weather) streamed via Starlink (Fig.
+2), enabling flexible validation from pure simulation to HIL testing with real measurements.
 TABLE II: Equipment Specifications
@@
 Inverter SolarEdge 3×10 kVA, single-phase
-Battery (TBI 2026) KORE Power 100 kW/100 kWh
+Battery (TBI 2026) KORE Power 100 kW/233 kWh
 LFP Battery Enphase ACB 1.0 6×1.2 kWh, 26.5 V
@@
 Regenerative Grid Simulator Chroma 61812 12 kVA, 0–600 V
-Microcontroller TI C2000 100 MHz
+Microcontroller TMDSCNCD28379D 100 MHz
 Real-Time Simulator Opal-RT OP5033 Intel® Xeon® (4 cores)
@@
 The ATL at the University of Vermont (Fig.
-2) integrates hardware and simulation capabilities to enable comprehensive testing of power electronics, energy storage, and control strategies in a reconfigurable cyber-physical environment.
-The bidirectional DC power supply feeds either the DC bus of two 5 kVA inverters connected in parallel or the batteries.
-The inverters, designed based on the UNIFI reference design [UNIFI2025], enable testing advanced control strategies and detailed investigation of inverter dynamics and their impact on the grid.
-Each inverter is controlled by a Texas Instruments (TI) TMDSCNCD28379D microcontroller, which enables the implementation of advanced control strategies by sensing analog voltage and current signals through ADC cards and generating required PWM signals based on control modes such as grid-forming (GFM) and grid-following (GFL) control, active/reactive power regulation, and coordinated multi-inverter operation.
+1) integrates hardware and simulation capabilities to enable comprehensive testing of power electronics, energy storage, and control strategies in a reconfigurable cyber-physical environment.
+A bidirectional DC power supply feeds either two parallel-connected 5 kVA inverters or the batteries.
+The inverters, designed based on the UNIFI reference design [2], enable testing of advanced control strategies and the investigation of inverter dynamics and their impact on the grid.
+Each inverter is controlled by a Texas Instruments (TI) microcontroller, which enables the implementation of control strategies such as grid-forming (GFM) and grid-following (GFL) control, active/reactive power regulation, and coordinated multi-inverter operation.
 In addition, the DC power supply is connected to six 1.2 kWh Enphase LFP batteries, allowing for systematic charge, discharge, and storage co-optimization studies.
 On the AC side, the inverters are connected to a regenerative grid simulator (RGS).
-The testbed is coupled to a 15 kW electrolyzer stack, which represents hydrogen production as a controllable load interfaced with the grid.
+The testbed is coupled to a 15 kW electrolyzer, which represents hydrogen production as a controllable grid-interfaced load.
 Furthermore, the Opal-RT simulator, with the OP4610 I/O expansion unit, provides the computational environment to emulate detailed grid behaviors.
 As shown in Fig.
-2, the testbed architecture is explicitly designed for ”plug-and-play” flexibility.
+1, the testbed architecture is explicitly designed for plug-and-play flexibility.
 This allows any component initially simulated within the Opal-RT to be seamlessly replaced by its physical hardware counterpart for full HIL validation.
@@
 II-B Hybrid Solar Test Center
-The ATL is supported by the HSTC located at the McNeil generating plant roughly one mile from campus.
-The HSTC hosts a 23 kW grid-connected solar PV array interfaced through SolarEdge inverters, along with an on-site weather station for real-time meteorological data acquisition.
-The weather station records solar irradiance, temperature, wind speed, and humidity every 5 seconds, providing synchronized environmental and electrical data for PV modeling and performance analysis.
-Future expansion includes the installation of a 100 kW/100 kWh battery energy storage system (BESS) in 2026, enabling field validation of battery-solar coordination strategies and grid services under real-world conditions.
-Figure 3: Data and communication architecture between HSTC and ATL.
-Environmental and operational data are collected from local and remote sources, as shown in Fig.
-3, to support real-time monitoring, control, and performance evaluation of all HES subsystems within the ATL.
-At the inverter and PV panel level, data from HSTC including AC and DC voltage, current, power, and frequency are measured using high-precision sensors and transmitted to the server through Modbus energy meter coupled with a communication gateway via Starlink network.
-All data streams are time-synchronized and stored on a dedicated server for visualization, analysis, and archival purposes.
-The setup provides a unique and unified platform for validating HES control strategies and system models, as well as assessing system dynamic behavior and grid-integration performance under controlled conditions.
-The key capabilities of the testbed are discussed in the following section.
+The ATL is supported by the HSTC located at the McNeil generating plant, one mile from campus.
+The HSTC hosts a 23 kW grid-connected solar PV array with SolarEdge inverters and an on-site weather station that records solar irradiance, temperature, wind speed, and humidity every 5 seconds, providing synchronized environmental and electrical data for PV modeling and performance analysis.
+A battery energy storage system (BESS) will be installed to enable field validation of battery-solar coordination strategies and grid services under real-world conditions.
+II-C Multi-site Communication and Data Integrity
+The ATL and HSTC facilities are interconnected through a communication architecture that employs a Starlink satellite network for real-time data streaming and an on-campus server.
+As shown in Fig.
+2, environmental and operational data from HSTC are measured using high-precision sensors and transmitted to a centralized server through Modbus coupled with a communication gateway.
+All data streams are time-synchronized and stored on the server for visualization and analysis, enabling real-time monitoring, control, and performance evaluation of HES subsystems.
+While multi-site architectures offer significant advantages for HES research by integrating realistic field data with laboratory HIL capabilities, they also presents technical challenges that must be carefully managed.
+Communication latency between geographically separated sites can degrade closed-loop control performance, especially for fast-responding applications such as frequency regulation and voltage control, which require sub-second coordination.
+Precise time synchronization across sites is critical to maintain data integrity.
+Furthermore, integrating diverse components with different communication protocols (Modbus, CAN Bus) creates interoperability complexity.
+High-frequency measurements from distributed sensors also generate large data volumes that require robust infrastructure for time-aligned storage, real-time processing, and post-experiment analysis.
+The architecture in this paper addresses these challenges through a robust centralized server for high-volume data management, protocol-specific scripts with built-in time synchronization for each communication standard, and direct data streaming via Starlink to reduce communication delays in control loops.
+Figure 2: Data-communication architecture for HSTC and ATL.
 III Platform Capabilities and Features
@@
 The ATL provides comprehensive experimental capabilities to validate physics-based and data-driven models across multiple components and timescales.
-The bidirectional DC power supply enables experimental validation of battery models through controlled charge-discharge cycling, including the energy reservoir model [omidi2025towards] and electro-thermal equivalent circuit model [mishra2024predictive] developed in prior work.
-For the electrolyzer subsystem, physics-based models describing hydrogen production rate, efficiency, and thermal dynamics have been developed and experimentally validated through systematic power cycling experiments on the 15 kW alkaline stack [ayubirad2025modeling].
-The platform further enables validation of efficiency-based and detailed switching inverter models.
-Data from the HSTC support validation of PV models across diverse operating conditions and temperature dependencies.
-At the system level, aggregated circuit models capture interactions between PV, battery, and inverter systems through equivalent circuit modeling frameworks, facilitating consistent state estimation across component boundaries and robust parameter identification [sang2024circuit].
+The bidirectional DC power supply enables validation of battery models through controlled charge-discharge cycling, including the energy reservoir model [12] and electro-thermal equivalent circuit model [10] developed in prior work.
+Physics-based electrolyzer models describing hydrogen production rate, efficiency, and thermal dynamics have been validated through power cycling experiments on the 15 kW alkaline stack [1].
+The platform also enables validation of efficiency-based and detailed switching inverter models.
+Data from the HSTC support PV model validation across diverse operating conditions and temperature dependencies.
+At the system level, aggregated circuit models capture interactions between PV, battery, and inverter systems through equivalent circuit modeling frameworks, facilitating state estimation and parameter identification [13].
 These validated models form the foundation for multi-objective optimization and control studies.
 III-B Control implementation and verification
-HES change how controllers must be designed.
-While individual components can often be controlled independently, the strong electrical and thermal coupling in HES makes coordinated controller design both necessary and beneficial.
+HES require coordinated controller design due to strong electrical and thermal coupling between components.
 As power systems become increasingly inverter dominated, traditional sources of inertia and voltage support disappear, and these services must be provided through tightly integrated, multi-layer control across HES components.
 Addressing these challenges requires experimental validation of advanced control and coordination strategies under realistic operating conditions.
-The ATL testbed bridges this gap by providing a flexible platform to validate the control algorithms at both the component and system levels.
-At the component level, ATL allows direct implementation and testing of inverter and battery control algorithms, such as GFM and GFL modes, active/reactive power regulation, Volt-VAR control, and multi-inverter coordination, under programmable conditions including voltage sags, weak grid operation, and fault events.
+The ATL testbed bridges this gap by providing a flexible platform to validate control algorithms at component and system levels.
+At the component level, ATL enables direct implementation and testing of inverter and battery control algorithms, such as GFM and GFL modes, active/reactive power regulation, Volt-VAR control, and multi-inverter coordination, under programmable conditions including voltage sags, weak grid operation, and fault events.
 At the system level, the platform supports hierarchical and distributed control schemes, including model predictive control and economic dispatch-based coordination of batteries, PV, and electrolyzers.
-These validated control strategies enable the ATL to test algorithms for providing different grid services, summarized in Table III.
-The shaded grid service is experimentally validated in the next section.
-TABLE III: Grid Services Capabilities of Campus Testbed
+These validated control strategies enable the ATL to test algorithms for providing grid services, summarized in Table III.
+The shaded grid service is experimentally validated in this paper.
+TABLE III: Capabilities of Multi-site Testbed
 Grid Services Time scale Participating assets
-Fast Frequency Response <<1s Battery+Inverter
+Synthetic damping/inertia <<1s Battery+Inverter
+Voltage regulation <<60s Battery/PV + Inverter
 Frequency regulation seconds-minutes Battery/Coupled HES
 […]
-Volt-Var and voltage regulation 100ms-10s Battery/PV + Inverter
-Power Quality <<1s Inverter
-Energy shifting and peak shaving minutes-hours Coupled HES
+Energy arbitrage & peak shaving minutes-hours Coupled HES
 III-C Hardware-in-the-loop (HIL) integration
@@
 It enables realistic evaluation of controllers, converters, and system interactions while maintaining safety and flexibility.
-The ATL testbed, equipped with an OPAL-RT real-time simulator and I/O unit, supports both CHIL and HIL.
-The OPAL-RT simulator supports steady-state, transient, and dynamic grid models, enabling comprehensive evaluation of hardware interactions with the grid across different timescales.
+The ATL testbed, equipped with an OPAL-RT real-time simulator and I/O unit, supports HIL configurations.
+The OPAL-RT simulator supports steady-state and dynamic grid models, enabling comprehensive evaluation of hardware interactions with the grid across different timescales.
 This enables detailed studies of inverter–grid interactions, coordination among HES components, and validation of advanced control strategies under realistic operating scenarios.
@@
 By coordinating PV generation with BESS operation, HES can smooth the power delivered to the grid.
-Grid operators impose ramp-rate limits to keep fluctuations within acceptable ranges, defined as:
+Grid operators impose ramp rate limits to keep fluctuations within acceptable ranges, defined as:
 $\text{Ramp Rate: }RR(t)=\frac{P_{\text{PV}}(t)-P_{\text{PV}}(t-\Delta t)}{\Delta t\cdot P_{\text{rated}}}\times 100$ (1)
 where $P_{\text{PV}}(t)$ is the PV power at time $t$, $\Delta t$ is the time interval, and $P_{\text{rated}}$ is the rated PV capacity.
-In this study, the maximum allowable ramp rate is set to $\pm 5\%$/min, which lies within the 2-10%/min range specified in several PV grid codes studies.
-To evaluate the smoothing capability at ATL, a CHIL experiment was conducted using the experimental setup shown in Fig.
-4.
-The experiment uses historical 5-second PV data from one inverter at the HSTC site.
-The PV profile is scaled to match the available battery capacity; this scaling does not affect ramp-rate behavior because ramp rates are expressed relative to the rated capacity.
-The OPAL-RT sends PV power and measured battery voltage to the TI controller, which runs a 30-minute moving-average smoothing algorithm, detailed in Algorithm 1.
-The controller issues battery current setpoints to the power supply through OPAL-RT.
-The battery management system (BMS) estimates the battery state of charge (SOC).
-Figure 4: Experimental setup for solar smoothing CHIL validation showing two series-connected Enphase batteries with a BMS, bidirectional DC power supply, TI control card, and OPAL-RT real-time simulator.
-In : $P_{\mathrm{PV}}[k]$, $V_{\mathrm{batt}}[k]$
-Out : $I_{\mathrm{set}}[k]$
+In this study, the maximum allowable ramp rate is set to $\pm 10~\%$/min, consistent with limits specified in PV grid codes [9].
+Figure 3: Experimental setup for solar smoothing HIL validation showing two series-connected Enphase batteries with a BMS, bidirectional DC power supply, TI control card, and OPAL-RT real-time simulator.
+Input : $P_{\mathrm{PV}}[k]$
+1exOutput : $P_{\mathrm{batt,h}}[k],\ P_{\mathrm{batt,s}}[k]$
+Parameters: $N,\ T_{s}$, $w_{h}\leftarrow\frac{P_{h}^{\max}}{P_{h}^{\max}+P_{s}^{\max}}$; $w_{s}\leftarrow 1-w_{h}$
+$P_{\mathrm{buf}}[i]\leftarrow 0,\ i=1{:}N$; $k\leftarrow 1$
 // Initialization
-$P_{\mathrm{buf}}[i]\leftarrow 0,\ i=1{:}N$; $k\leftarrow 1$
 while controller is running do
+ 1ex
+ $P_{\mathrm{buf}}[([k-1]\bmod N)+1]\leftarrow P_{\mathrm{PV}}[k]$
   // 1) Insert new PV sample
-  $P_{\mathrm{buf}}[([k-1]\bmod N)+1]\leftarrow P_{\mathrm{PV}}[k]$
-  // 2) Compute smoothed power
  $\hat{P}_{\mathrm{PV}}[k]\leftarrow\frac{1}{N}\sum_{i=1}^{N}P_{\mathrm{buf}}[i]$
-  // 3) Battery power
+  // 2) Smoothed PV power
  $P_{\mathrm{batt}}[k]\leftarrow P_{\mathrm{PV}}[k]-\hat{P}_{\mathrm{PV}}[k]$
-  // 4) Current setpoint
-  $I_{\mathrm{set}}[k]\leftarrow\dfrac{P_{\mathrm{batt}}[k]}{V_{\mathrm{batt}}[k]}$
-  // 5) Output and increment
-  Output $I_{\mathrm{set}}[k]$
-  $k\leftarrow k+1$ ; sleep($T_{s}$)
+  // 3) Total battery command
+ $P_{\mathrm{batt,h}}[k]\leftarrow w_{h}\cdot P_{\mathrm{batt}}[k]$
+  // 4) Battery power allocation
+  $P_{\mathrm{batt,s}}[k]\leftarrow w_{s}\cdot P_{\mathrm{batt}}[k]$
+ Output $P_{\mathrm{batt,h}}[k],\ P_{\mathrm{batt,s}}[k]$
+  // 5) Output
+  $k\leftarrow k+1$, sleep($T_{s}$)
 Algorithm 1 PV Smoothing Controller
-A two-hour snapshot of the CHIL experiment, shown in Fig.
-5, highlights the system smoothing performance.
+To evaluate the smoothing capability at ATL, an HIL experiment was conducted using the experimental setup shown in Fig.
+3.
+The experiment uses historical 5-second PV power data from the HSTC site.
+To meet the power requirements for PV smoothing, the testbed integrates physical battery hardware (1.2 kW/2.4 kWh) together with three battery equivalent circuit models connected in series and implemented as the simulated battery (7.2 kW/7.2 kWh).
+The OPAL-RT sends PV power to the TI controller, which runs a 30-minute moving-average smoothing algorithm, detailed in Algorithm 1.
+An energy management system (EMS) distributes the required battery power between the two battery systems based on their maximum power ratings.
+The controller issues setpoints for the battery to the power supply through OPAL-RT.
+The battery management system (BMS) estimates the battery SOC.
 Fig.
-5(a) shows the raw PV power and compares the smoothed results from simulation and experiment, which indicate that the results align closely.
+4 presents a two-hour snapshot of the HIL experiment.
+The raw PV power exhibits a maximum ramp rate of 56 %/min, significantly exceeding the $\pm 10~\%$/min grid limit.
+Fig.4(a) demonstrates that the BESS effectively reduces this to 9.45%/min.
+The simulation and experimental results show close agreement, with a root mean square error (RMSE) of 0.25 kW (2.4 % of maximum power).
+This error is attributed to communication latency, ADC quantization, sensor noise, and sampling effects.
 Fig.
-5(b) presents the dispatched battery power, and Fig.
-5(c) shows the resulting SOC, which remain within safe operating limits throughout the experiment.
-The raw PV profile shows a maximum ramp rate of 56%/min, exceeding the $\pm 5\%$/min grid limit, whereas the BESS effectively reduces the maximum ramp rate to 3.75%/min.
-The corresponding distribution of ramp rates is shown in Fig.
-5(d), which illustrates the significant reduction in extreme events.
-These results validate the capability of the ATL facility to effectively implement and test PV smoothing strategies using BESS.
-Figure 5: Two-hour snapshot of HIL experimental PV power smoothing using historical 5-second HSTC data and ATL battery setup.
-(a) raw and smoothed PV power, here, $P_{\text{PV}}$ denotes the raw PV power, $\hat{P}_{\text{PV}}^{e}$ and $\hat{P}_{\text{PV}}^{s}$ denote the smoothed PV power from experiment and simulation, respectively, (b) dispatched battery power, (c) battery SOC, and (d) distribution of ramp rates.
+4(b) presents dispatched battery power for the physical and simulated batteries.
+Fig.
+4(c) shows the SOC of both batteries.
+The physical battery exhibits larger deviations due to its lower capacity and maximum power; however, both batteries remain within safe operating limits.
+Fig.
+4(d) illustrates the ramp rate distribution, showing reduced extreme events and zero grid-constraint violations.
+These results validate ATL’s capability to implement and test PV smoothing strategies under multi-site operation.
+Figure 4: Two-hour snapshot of HIL experimental PV power smoothing using 5-second HSTC data and the ATL battery setup.
+(a) Raw and smoothed PV power, where $P_{\text{PV}}$ denotes the raw PV power, and $\hat{P}_{\text{PV}}^{e}$ and $\hat{P}_{\text{PV}}^{s}$ denote the experimentally measured and simulated smoothed PV power, respectively, (b) dispatched battery power of the physical and simulated batteries, (c) battery SOC, and (d) ramp-rate distributions.
 V Conclusion